OpenStudy (anonymous):

The price of a car has been reduced from $16,000 to $11,200. What is the percentage decrease of the price of the car? a 3% b 48% c 30% d 70%

OpenStudy (anonymous):

@lilshane

OpenStudy (anonymous):

@Inquus

OpenStudy (anonymous):

@jagr2713

OpenStudy (anonymous):

@Mr_Perfection_xD

OpenStudy (anonymous):

\[(\frac{ current }{ original })* 100 = \] (percent of original) 100-(percent of original) = percent change so \[(\frac{ 11200 }{ 16000})*100 = 70\] \[100 - 70 = 30\] change of price (in percent) is 30%

OpenStudy (anonymous):

C 30%

OpenStudy (anonymous):

C 30%

OpenStudy (anonymous):

thanks your awesome

OpenStudy (anonymous):

No problem, you're welcome!